For the given equations, find a) f(0) b) the zeros of f
(I got the first part, a, but I didn't understand the second part, b. It would be great if someone helped :) )
1. f(x) = 4x - 20
2. f(x) = x^2 + 4x - 12

Answer by lwsshak3(11628)

For the given equations, find a) f(0) b) the zeros of f
(I got the first part, a, but I didn't understand the second part, b. It would be great if someone helped :) )
1. f(x) = 4x - 20
a) f(0)=0-20=-20
b) set f(x)=0
0=4x-20
4x=20
x=5 (a zero of f)
2. f(x) = x^2 + 4x - 12
a) f(0)=0+0-12=-12
b) set f(x)=0
0=x^2+4x-12
x^2+4x-12=0
(x+6)(x-2)=0
x+6=0
x=-6
or
x-2=0
x=2
zeros of f are -6 and 2